# Efficient Optimization Accelerator Framework for Multi-state Ising Problems

### v1

@@ Abstract
Ising Machines are a prominent class of hardware architectures that aim to solve NP-Hard combinatorial optimization problems. These machines consist of a network of interacting binary spins/neurons that evolve to represent the optimum ground state energy solution. Generally, combinatorial problems are transformed into quadratic unconstrained binary optimization (QUBO) form to harness the computational efficiency of these Ising machines. However, this transformation, especially for multi-state problems, often leads to a more complex exploration landscape than the original problem, thus severely impacting the solution quality. To address this challenge, we model the spin interactions as a generalized boolean logic function to significantly reduce the exploration space. We benchmark the graph coloring problem from the class of multi-state NP-hard optimization using probabilistic Ising solvers to illustrate the effectiveness of our framework. The proposed methodology achieves similar accuracy compared to state-of-the-art heuristics and machine learning algorithms, and demonstrates significant improvement over the existing Ising methods. Additionally, we demonstrate that combining parallel tempering with our existing framework further reduces the coloring error by up to 50 compared to the conventionally used Gibbs sampling algorithm. We also design a 1024-neuron all-to-all connected probabilistic Ising accelerator that shows up to 10000 performance acceleration compared to heuristics while reducing the number of required physical neurons by 1.5-4 compared to conventional Ising machines. Indeed, this accelerator solution demonstrates improvement across all metrics over the current methods, i.e., energy, performance, area, and solution quality. Thus, this work expands the potential of existing Ising hardware to solve a broad class of these multistate optimization problems.
@@
Department of Electrical Engineering and Computer Sciences, University of California, Berkeley, California 94720, USA
@@
New computing paradigms are getting significant attention due to exponentially growing computing needs [1]. A wide variety of problems falls into the class of non-deterministic polynomial-time hard (NP-hard) and are difficult to solve optimally using conventional computing solutions [2, 3]. The solution space grows exponentially with problem size, therefore brute force searching does not work. However, heuristics [4] and annealing [2] based approaches have been conventionally used to tackle these computationally hard problems in domains such as logistics [5, 6], biology [7], integrated circuits design [8], etc. In this context, Ising machines-based accelerators [9, 10, 11, 12, 13, 14, 15, 16] are currently being leveraged to efficiently find the solution to hard optimization problems. Various technologies and hardware architectures have been explored to build these Ising accelerators including quantum annealing with superconducting qubits [17], classical annealing in memristor or RRAM [18, 19], coherent Ising machines employing optical oscillators [13, 20], coupled oscillators [14, 16, 21], neuromorphic hardware [22] and stochastic circuits (probabilistic bits) [12, 9, 23, 10]. Quantum annealers show promising results [24] but require low temperatures making them expensive in cost and power. Therefore, classical alternatives have gained attention due to their room-temperature operation and realization using the current semiconductor process flow. This work particularly focuses on stochastic/probabilistic Ising machines to efficiently solve combinatorial optimization problems.
@@
Graph coloring is an NP-hard optimization problem that seeks to assign different colors to the connected nodes of a graph network. It is an example of an integer optimization problem where each integral value represents the color of a node. Previously, annealing and heuristic methods (Tabu-search [32]) were utilized to tackle this problem. These approaches achieve good solution accuracy but suffer from long runtimes for large and densely connected problem instances. Recently, learning-based approaches[33, 34] especially graph neural networks are applied to solve the problem accurately and efficiently at a scale. Conventional Ising hardware or its modified versions are also proposed to solve this problem. However, they significantly lag behind heuristics and learning-based approaches in solution quality. To an extent, earlier works [18, 30, 31] overcome this bottleneck by adopting post-processing techniques using additional hardware and software, which adversely affects area and computation time.
@@
In this work, we present an end-to-end probabilistic Ising implementation that combines advances in multi-state problem mapping, spin interaction design, and an efficient hardware architecture to demonstrate significant improvement in area, solution accuracy, and time-to-solution. First, we propose vectorized mapping that represents the node colors as binary vectors rather than using the customary one-hot form. This circumvents the additional mapping constraint in the Ising Hamiltonian that arises due to one-hot encoding [27]. Thus, it completely discards the exploration of infeasible/invalid solution space and improves the solution quality. Second, the interactions among the binary vector states are modeled using truth tables employed in digital logic. Third, we implement an efficient FPGA accelerator that uses higher-order multiplexers to represent these truth tables. Altogether, this multi-state Ising implementation shows approximately 10000 speed improvement and 5 power improvement compared to GPUs for solving problems up to 256 nodes and 16 colors graph coloring problems. This current methodology is then integrated with the parallel tempering approach to improve the solution quality and provide competitive or even better coloring results than Tabucol heuristics and machine learning algorithms.
Mapping Framework
@@
A wide range of combinatorial optimization problems belongs to multistate/integer-valued solution space, including graph coloring, knapsack problems [35], etc. Graph coloring problem, investigated in this work, has applications, such as layout decomposition[36], register allocation[37], logic minimization[38], scheduling[33], and many more. The problem instance that aims to color nodes with colors, requires physical neurons/nodes in the conventional Ising framework. Because this framework relies on one-hot encoding to represent node colors, additional constraints are imposed in Hamiltonian to enforce the legal combination of one-hot bits [27]. It is written as:
@@
where represents one-hot bit for node . This soft constraint Hamiltonian is expected to be zero for the ground state energy solution. However, it often takes the non-zero value, thereby making Ising samplers inefficient in solving such problems. (see Fig. 2a). The one-hot encoding requires bits to represent possible colors of a node. As a result, Ising machines search over states with only valid states in the solution space. It makes the optimization hard and deteriorates the solution quality. Hence, the conventional Ising framework often fails to solve this class of problems. Despite these challenges, this approach is often adopted for most multistate optimization problems because it leads to quadratic interactions supported in most Ising machines [26, 10, 16, 20]. In probabilistic computing [12], binary multiplexers offer simplistic and efficient realization for state-weight multiplication.
@@
In this work, we propose a vectorized mapping approach (Fig. 2b) to tailor the graph coloring problem using a binary encoding technique. It represents the color state using bits vector and requires physical neurons/nodes for a graph coloring problem with nodes and colors. In this way, it eliminates all invalid state space and removes the extra constraints in Ising Hamiltonian (Eq. 1). Therefore, the proposed approach leads to a solvable energy landscape without any additional complexities. We also adopt parallel tempering [39] to enhance the solution exploration and improve the solution quality.
@@
The proposed vector mapping is a generic approach that can be applied to any multistate optimization problem being solved on Ising accelerators (see Fig. 2b). First, we model the function (see Fig. 2b) in the Hamiltonian for a specific problem. Here multiplication factor enforces edge connectivity. is modeled in a truth table-based format. These truth tables can then be represented via multiplexers to develop a probabilistic accelerator employing the proposed mapping. For graph coloring problems, for the cases when the nodes ( and ) take the same value or color value greater than q (invalid colors), becomes one. is zero for the rest of the cases. The resulting Hamiltonian is a higher-order polynomial. It is minimized when is zero, implying graph coloring criteria are satisfied. The hardware architecture for this scheme is derived from the standard node update rule for probabilistic Ising machines. For bit of node , the update rule is as follows:
@@
where takes binary value 0 and 1. The binary nature of converts the differentiation transforms into equal to - . This is implemented in hardware using higher-order multiplexers followed by a subtraction unit (see Fig. 2), called as vecmul unit. The rest of the hardware implementation remains the same as shown in Fig. 1.
Results
@@
Next we investigate the proposed vectorized mapping approach for the graph coloring problem. For benchmarking, we have used it on publicly available COLOR dataset [40]. First, we implement both probabilistic Ising and proposed vectorized approaches on GPU for comparison. Fig. 3 shows the results for the problem, which is deemed a hard problem [33]. For both schemes, the exploration for solution happens in a similar way (Fig. 3a). For this particular instance, none of the states explored by the Ising framework is valid and therefore unable to find reasonable solutions for most of the problem instances (Fig. 3b) (for more details, see Supplementary section S1). This comparison for other problem instances in the dataset is shown in the Extended Data Fig. S1 and S2. Further, we employ the parallel tempering method to enhance the exploration while sampling the energy landscape. This method involves the simultaneous simulation of M replica systems with vectorized mapping running at different temperature schedules. At high temperatures, the system tends to explore the broad region of the solution space while low temperatures allow precise sampling around a particular region (Fig. 2b). High-temperature replicas ensure that the system doesn’t get stuck around the local minimum solution. The states of higher-temperature replica systems are exchanged with low-temperature systems via a swap rule (see Supplementary Section 2). This state exchange phenomenon enables the exploration of certain parts of solution space which could not be possible via a single temperature schedule. Fig. 3c depicts the energy exploration for the problem using the parallel tempering approach. It clearly shows that the system continuously escapes from the minimum local energy solution and therefore, achieves the best possible solution accuracy.
@@
We benchmark the vectorized mapping solution accuracy for graph coloring problems in the dataset against learning-based approaches including graph neural network (GNNs) [34] and its derivative architectures GraphSage (PI-SAGE) [33] and heuristics-based approach, i.e., Tabucol [32]. GNN-based solvers frame the coloring problem as a multi-class node classification problem and employ unsupervised training by formulating a loss function. In contrast, the Tabucol technique searches the ground state energy solution by moving small steps and maintains a tabu list to avoid cycling around local minima. We implement Tabucol heuristics on Intel Xeon Gold 6330 processor and solve the graph coloring problem instances 200 times with 1000 iteration steps each. For accuracy comparison, we also run Ising and vectorized mapping framework with single-flip Gibbs sampling on Nvidia A100 Tensor Core GPU and solve the same problems 200 times with 1000 iteration steps each. The results comparing the described methods are shown in Table 1. It includes easy, medium, and hard problem instances labeled in work [41]. The Ising approach performs worse by only being able to solve small and easy problem instances. Among GNN-based solvers, the GraphSage architecture (PI-SAGE) offers better solution accuracy but it suffers from longer training times [33]. By contrast, proposed vectorized mapping gives competitive coloring results compared to Tabucol heuristics and PI-SAGE GNN while having slightly lower accuracy for hard-to-solve problem instances. Employing the described parallel tempering with vectorized mapping reduces the error up to 50% on these hard problems and therefore performs better than other methods.
@@
We implement the proposed vectorized architecture (Fig. 2) onto VCU118 FPGA to establish the performance acceleration and energy efficiency benchmarks. The FPGA accelerator uses the memory-mapped IO interface used by software applications to program the problem weights and take the solution out (see Supplementary Section 3 for more details). The accelerator (Fig. 1) fetches the weight data from the memory and does neuron states-weight multiplication. The accumulated product is passed through a sigmoid activation implemented using LUT and further compared with a Linear Feedback Shift Register (LFSR) based random number generator (RNG) to get the updated node value. The architecture runs at 90 MHz clock frequency where only one node gets updated in each cycle. Further, it supports graph coloring problem sizes up to 256 nodes all-to-all connected with a maximum chromatic number of 16, equivalent to 1024 vectorized nodes in total. The framework achieves the same accuracy for dataset problems as the GPU-based vectorized mapping (see Table 1).
@@
The proposed architecture not only gives better accuracy than the baseline Ising-based architecture but also reduces the area implementation. The area implementation is quantified in terms of number of physical nodes required to map any problem instance and solve it. The increase in physical nodes leads to an increase in the number of interactions and hence accumulator size grows leading to a penalty on the implementation area. Fig. 4a shows that the vectorized mapping requires 1.5-4 times fewer nodes for the graph coloring problem instances with chromatic numbers up to 16.
@@
The FPGA implementation of the proposed approach shows 1000 - 10000 speedup compared to Tabucol heuristics implemented on Intel Xeon Gold 6330 CPU cores and its equivalent implementation on Nvidia A100 Tensor Core GPU (see Fig. 4b). Vectorized mapping on GPU could not take advantage of efficient multiplication of binary states with weights and, therefore, suffers with time-to-solution. The same hypothesis will hold for the accelerators that are not built to support these binary calculations. The FPGA accelerator also offers 5 power improvements over GPU implementation. Therefore, accelerating the time-to-solution by 3-4 orders of magnitude on FPGA at a low power budget significantly improves energy efficiency.
Conclusion

### v2

# Efficient Optimization Accelerator Framework for Multi-state Spin Ising Problems

@@ Abstract
Ising Machines are emerging hardware architectures that efficiently solve NP-Hard combinatorial optimization problems. Generally, combinatorial problems are transformed into quadratic unconstrained binary optimization (QUBO) form, but this transformation often complicates the solution landscape, degrading performance, especially for multi-state problems. To address this challenge, we model spin interactions as generalized boolean logic function to significantly reduce the exploration space. We demonstrate the effectiveness of our approach on graph coloring problem using probabilistic Ising solvers, achieving similar accuracy compared to state-of-the-art heuristics and machine learning algorithms. It also shows significant improvement over state-of-the-art QUBO-based Ising solvers, including probabilistic Ising and simulated bifurcation machines. We also design 1024-neuron all-to-all connected probabilistic Ising accelerator on FPGA with the proposed approach that shows 10000 performance acceleration compared to GPU-based Tabucol heuristics and reducing physical neurons by 1.5–4 over baseline Ising frameworks. Thus, this work establishes superior efficiency, scalability and solution quality for multi-state optimization problems.
@@
Department of Electrical Engineering and Computer Sciences, University of California, Berkeley, CA 94720, USA
[…]
Introduction
@@
New computing paradigms are getting significant attention due to exponentially growing computing needs [1]. A wide variety of problems falls into the class of non-deterministic polynomial-time hard (NP-hard) and are difficult to solve optimally using conventional computing solutions [2, 3]. The solution space grows exponentially with problem size, therefore making brute force searching impractical for large problem instances. However, heuristics [4] and annealing [2] based approaches have been conventionally used to tackle these computationally hard problems in domains such as logistics [5, 6], biology [7], integrated circuits design [8], etc. In this context, Ising machines-based accelerators [9, 10, 11, 12, 13, 14, 15, 16] are currently being leveraged to efficiently find the solution to hard optimization problems. Various technologies and hardware architectures have been explored to build these Ising accelerators including quantum annealing with superconducting qubits [17], classical annealing in memristor or RRAM [18, 19], coherent Ising machines employing optical oscillators [13, 20], coupled oscillators [14, 16, 21], neuromorphic hardware [22] and stochastic circuits (probabilistic bits) [12, 9, 23, 10]. Quantum annealers show promising results [24] but require low temperatures making them expensive in cost and power. Therefore, classical alternatives have gained attention due to their room-temperature operation and realization using the current semiconductor process flow. This work particularly focuses on stochastic/probabilistic Ising machines to efficiently solve combinatorial optimization problems.
@@
Graph coloring is an NP-hard optimization problem that seeks to assign different colors to the connected nodes of a graph network. It is an example of an integer optimization problem where each integral value represents the color of a node. Previously, annealing and heuristic methods (Tabu-search [32]) were utilized to tackle this problem. These approaches achieve good solution accuracy but suffer from long runtimes for large and densely connected problem instances. Recently, learning-based approaches[33, 34] especially graph neural networks are applied to solve the problem accurately and efficiently at a scale. Existing Ising hardware or its modified versions are also proposed to solve this problem. However, they significantly lag behind heuristics and learning-based approaches in solution quality. To an extent, earlier works [18, 30, 31] overcome this bottleneck by adopting post-processing techniques using additional hardware and software, which adversely affects area and computation time.
@@
In this work, we present an end-to-end probabilistic Ising implementation that combines advances in multi-state problem mapping, spin interaction design, and an efficient hardware architecture to demonstrate significant improvement in area, solution accuracy, and time-to-solution. First, we propose vectorized mapping that represents the node colors as binary vectors rather than using the customary one-hot form. This circumvents the additional mapping constraint in the Ising Hamiltonian that arises due to one-hot encoding [27]. Thus, it completely discards the exploration of infeasible/invalid solution space and improves the solution quality. Second, the interactions among the binary vector states are modeled using truth tables employed in digital logic. Third, we implement an efficient FPGA accelerator that uses higher-order multiplexers to represent these truth tables. Altogether, this multi-state Ising implementation shows approximately 100000 speed improvement and 5 power improvement compared to its GPU-based implementation for solving problems up to 256 nodes and 16 colors graph coloring problems. This current methodology is then integrated with the parallel tempering approach to improve the solution quality and provide competitive or even better coloring results than Tabucol heuristics and machine learning algorithms.
1 Ising Mapping Framework
@@
A wide range of combinatorial optimization problems belongs to multistate/integer-valued solution space, including graph coloring, knapsack problems [35], etc. Graph coloring problem, investigated in this work, has applications, such as layout decomposition[36], register allocation[37], logic minimization[38], scheduling[33], and many more. The problem instance that aims to color nodes with colors, requires physical neurons/nodes in the Ising framework. Because this framework relies on one-hot encoding to represent node colors, additional constraints are imposed in problem Hamiltonian ( in Eq. 1) to enforce the legal combination of one-hot bits [27] (represented by Eq. 2). Therefore, the Ising Hamiltonian (H) for graph coloring problem becomes .
@@
In the equations, represents one-hot bit for node , denotes the set of all edges in the problem graph, is the connectivity weighting factor, and is the one-hot constraint factor. This soft constraint Hamiltonian is expected to be zero for the ground state energy solution. However, it often takes the non-zero value, thereby making Ising samplers inefficient in solving such problems. (see Fig. 2a). The one-hot encoding requires bits to represent possible colors of a node. As a result, Ising machines search over states with only valid states in the solution space. It makes the optimization hard and deteriorates the solution quality. Hence, the Ising frameworks following one-hot encoding often fail to solve this class of problems. Despite these challenges, this approach is often adopted for most multistate optimization problems because it leads to quadratic interactions supported in most Ising machines [26, 10, 16, 20]. In probabilistic computing [12], binary multiplexers offer simplistic and efficient realization for state-weight multiplication.
2 Vectorized Mapping Framework
@@
In this work, we propose a vectorized mapping approach (Fig. 2b) to tailor the graph coloring problem using a binary encoding technique. It represents the color state using the () bit vector, which means color for node is represented as {}. Therefore, this mapping requires physical neurons/nodes for a graph coloring problem with nodes and colors. In this way, it eliminates all invalid state space and removes the extra constraints in Ising Hamiltonian (Eq. 2). Therefore, the proposed approach leads to a solvable energy landscape without additional complexities. We also adopt parallel tempering [39] to enhance the solution exploration and improve the solution quality.
@@
The proposed vector mapping is a generic approach that can be applied to any multistate optimization problem being solved on Ising accelerators (see Fig. 2b). This approach represents the states in binary representation and only requires modeling the function operator in the Hamiltonian H ( Eq. 3) for a specific problem. In Eq. 3 , represents the edge (E) weight connecting the nodes and , and function operator is modeled in a truth table-based format. Algorithm 1 describes the formulation of the truth table for the graph coloring problem and Fig. 2 shows the formulation of the truth table of a 4 nodes, 3 colors graph coloring problem. In the cases when nodes ( and ) take the same value or color value greater than q (invalid colors), becomes one. is zero for the rest of the cases. The resulting Hamiltonian is a higher-order polynomial due to the operator function. Ideally, graph coloring criteria are met when the Hamiltonian energy reaches its global minimum of zero, which corresponds to the condition where is equal to zero. Further, the truth table-based algorithm facilitates its representation as multiplexers, particularly in the case of Ising hardware, which digitally models spin interactions. In this work, we leverage the probabilistic Ising machines framework to develop the accelerator following the proposed vectorized mapping. The hardware architecture for this scheme is derived from the standard node update rule for probabilistic Ising machines. For bit of node , the update rule is as follows:
@@
where takes binary value 0 and 1, is temperature coefficient and is sigmoid function ( for input value ). The binary nature of converts the differentiation transforms into equal to - . This is implemented in hardware using higher-order multiplexers followed by a subtraction unit (see Fig. 2), called as vecmul unit. The rest of the hardware implementation remains the same as shown in Fig. 1.
3 Accuracy Analysis and Benchmarks
@@
Next we investigate the proposed vectorized mapping approach for the graph coloring problem. For benchmarking, we have used it on publicly available COLOR dataset [40]. First, we implement probabilistic Ising frameworks following one-hot encoding (Ising) and proposed binary encoding (vectorized) approaches on GPU for comparison. Fig. 3 shows the results for the problem, which is deemed a hard problem [33]. For both schemes, the exploration for solution happens in a similar way (Fig. 3a). Fig. 3b illustrates that the vectorized mapping gives superior coloring results as compared to Ising framework. Even with an optimal set of connectivity weight factor and one-hot constraint weight factor parameters (see Supplementary Fig. S1), the effectiveness of the Ising framework is limited by the one-hot encoding constraint in Eq.2. This comparison for other problem instances in the dataset is shown in the Supplementary Fig. S3 and S4. Further, we employ the parallel tempering method to enhance the exploration while sampling the energy landscape. This method involves the simultaneous simulation of M replica systems with vectorized mapping running at different temperature schedules. At high temperatures, the system tends to explore the broad region of the solution space while low temperatures allow precise sampling around a particular region (Fig. 2b). High-temperature replicas ensure that the system doesn’t get stuck around the local minimum solution. The states of higher-temperature replica systems are exchanged with low-temperature systems via a swap rule (see Methods Section 1). This state exchange phenomenon enables the exploration of certain parts of solution space which could not be possible via a single temperature schedule. Fig. 3c depicts the energy exploration for the problem using the parallel tempering approach. It clearly shows that the system continuously escapes from the minimum local energy solution and therefore, achieves the best possible solution accuracy.
@@
We evaluate the accuracy of vectorized mapping solutions against learning-based methods (graph neural network [34] and its GraphSage architecture PI-SAGE [33] ), heuristic approaches (Tabucol [32, 41]), and state-of-the-art Ising solvers (probabilistic Ising machines [42], simulated bifurcation[43, 44] ) for graph coloring across standard benchmark datasets. GNN-based solvers frame the coloring problem as a multi-class node classification problem and employ unsupervised training by formulating a loss function. In contrast, the Tabucol technique searches the ground state energy solution by moving small steps and maintains a tabu list to avoid cycling around local minima. We run Tabucol heuristics, probabilistic Ising machines, and vectorized mapping framework with single-flip Gibbs sampling on Nvidia A100 Tensor Core GPU and solve the same problems 200 times with 1000 iteration (update) steps each. Additionally, simulated bifurcation machines algorithm has been run on the same GPU solving the same problems 200 times with 10000 iteration steps each. We report best possible results achieved by the described methods and compares them in Table 1 and 2. It includes easy, medium, and hard problem instances labeled in work [45]. The Ising approach performs worse by only being able to solve small and easy problem instances accurately. Among GNN-based solvers, the GraphSage architecture (PI-SAGE) offers better solution accuracy but it suffers from longer training times [33]. By contrast, proposed vectorized mapping gives competitive coloring results compared to Tabucol heuristics and PI-SAGE GNN while having slightly lower accuracy for hard-to-solve problem instances. Employing the described parallel tempering with vectorized mapping reduces the error up to 50% on these hard problems and therefore performs better than other methods.
4 Evaluation of Success Probability and Time-to-Solution
We employ the success probability metric extensively used in other works[16, 46] that captures the solution quality of statistical algorithms. We define error as number of incorrectly colored edges divided by total edges in the problem graph. To calculate success probability (), we run each coloring problem for 200 times (or parallel runs) using each algorithm and calculate the success in getting the error less than 2 for those iterations. The results in Fig. 4b confirms that the vectorized mapping achieves competitive success probability against the Tabucol heuristic approach and produces a better quality solution compared to state-of-the-art Ising solvers, including probabilistic Ising machine and Simulated Bifurcation machines. Additionally, time-to-solution (TTS) in Eq. 5 is defined as the time needed to obtain a solution within a specified accuracy across multiple runs, with a probability of 99. represents the average time to complete a single run. The algorithms that achieve success probability greater than 99, TTS is defined as the average time to reach the solution across parallel runs. Using this methodology, Fig. 4c reports the TTS for different statistical algorithms used to solve the graph coloring problem instances. Overall, the success probability and TTS metric capture a critical aspect of solver performance. A higher success probability reflects a solver’s ability to tackle problem instances more efficiently, requiring fewer attempts. This efficiency is represented by incorporating a factor in TTS formulation that accounts for the influence of success probability, providing a robust evaluation of solver effectiveness in terms of both accuracy and computational efficiency.
5 FPGA Accelerator Implementation and Results
@@
We implement the proposed vectorized architecture (Fig. 2b) onto VCU118 FPGA to establish the performance acceleration and energy efficiency benchmarks. The FPGA accelerator uses the memory-mapped IO interface used by software applications to program the problem weights and take the solution out (see Methods Section 2 for more details). The accelerator (Fig. 1) fetches the weight data from the memory and does neuron states-weight multiplication. This process includes computing the F-operator, defined in Algorithm 1 which is implemented in hardware using truth tables or higher-order multiplexers, as illustrated in Fig. 2b. The structure of this operator is largely determined by the graph’s coloring constraints. The accumulated product, represented with 8-bit precision, is passed through a sigmoid activation function that is implemented using a look-up table (LUT) containing (256) entries. The LUT generates a 16-bit output, which is then compared with a 16-bit random number generated using a Linear Feedback Shift Register (LFSR) to get the updated node value. The architecture runs at 90 MHz clock frequency where only one node gets updated in each cycle. Further, it supports graph coloring problem sizes up to 256 nodes all-to-all connected with a maximum chromatic number of 16, equivalent to 1024 vectorized nodes in total. The framework achieves the same accuracy for dataset problems as the GPU-based vectorized mapping (see Table 1).
@@
The proposed architecture not only gives better accuracy than the baseline Ising-based architecture following one-hot encoding but also reduces the area implementation. The area implementation is quantified in terms of number of physical nodes required to map any problem instance and solve it. The increase in physical nodes leads to an increase in the number of interactions and hence accumulator size grows leading to a penalty on the implementation area. Fig. 4a shows that the vectorized mapping requires 1.5-4 times fewer nodes for the graph coloring problem instances with chromatic numbers up to 16.
The FPGA architecture also uses single flip-Gibbs sampling to update the node, therefore, N nodes need to be updated for vectorized mapping system for one complete iteration or time step. Thus, the time complexity for the proposed framework will scale with (). The time update for a single node update affects the prefactor of thes mentioned time complexity. Specifically, in probabilistic Ising/vectorized, this prefactor is contribution of multiplexer (states-weight multiplication), accumulator, sigmoid activation calculation and comparator delays. Owing to binary nature of nodes, the accumulator dominates the node update timing[9]. As a result, introducing higher-order multiplexers has minimal impact on the overall time complexity scaling.
@@
The FPGA implementation of the proposed vectorized mapping approach shows 10000 speedup compared to Tabucol heuristics and Ising solvers (Probabilistic Ising and Simulated Bifurcation) implementation on Nvidia A100 Tensor Core GPU (see Fig. 4c). Vectorized mapping on GPU could not take advantage of efficient multiplication of binary states with weights and, provides only comparable time-to-solution to heuristics. This same trend is expected to hold for any accelerators that are is built to support these binary calculations. Similarly, Ising mapping-based solvers, including Probabilistic Ising and Simulated Bifurcation, can benefit from FPGA acceleration [42, 47], potentially narrowing the performance gap with our FPGA implementation of the vectorized mapping. Nevertheless, for problems with more than 100 nodes, these solvers may continue to yield sub-optimal solutions. The FPGA accelerator also offers 5 power improvements over GPU-based vectorized mapping implementation. Therefore, accelerating the time-to-solution by > 4 orders of magnitude on FPGA at a low power budget significantly improves energy efficiency Fig. 4d.
One question that may arise is: is the acceleration achieved for the proposed method over Heuristics solely due to the fact that the vectorized method was implemented on FPGA? In this regard, we note here that the Heuristics take advantage of sequential strategies that do not scale well on parallel architecture. This has been studied extensively in the literature. For example, in Supplementary Fig. S10a we show a comparison between CPU and GPU implementations of Tabucol. It is clearly seen that the GPU implementation provides virtually no acceleration. By contrast GPU implementation of the proposed vectorized method shows large acceleration as shown in Supplementary Fig. S10b. This underscores a strength of the proposed method that makes it amenable for scaling on specialized hardware such as the FPGA.
Discussion